### v1

# Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star forming galaxies at z5.2 with the N2CLS

Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star forming galaxies at z5.2 with the N2CLS
High-density environments, such as early galaxy overdensities, play a critical role in hierarchical structure formation and galaxy evolution, providing an ideal setting for accelerated evolution of galaxies. As part of the NIKA2 Cosmological Legacy Survey, we report the discovery of a remarkable overdensity of ten bright dusty star-forming galaxies at z5.2 in the GOODS-N field. Among these, three galaxies with spectroscopic redshifts (N2GN_1_01, 06, and 23, known as GN10, HFD850.1, and S3, respectively) had already been identified as part of the exceptional overdensity at z5.1-5.3. This structure is notably elongated along the line-of-sight, spanning approximately 30 cMpc. We report the discovery of N2GN_1_13 at zspec=5.182, a massive dusty star-forming galaxy that we confirmed as part of the overdensity through targeted NOEMA follow-up observations of the N2CLS. In addition, by combining our analysis with JWST/FRESCO data, we identified another dusty galaxy at zspec=5.201 (N2GN_1_61). In addition to these five spectroscopically confirmed galaxies, we identify five more with photometric redshifts consistent with the overdense structure. These galaxies are massive (with a median stellar mass of 1011 M⊙) and highly obscured (with a median AV of 2.9), caught in a short-lived yet extreme starburst phase at z5.2. Their high star formation rates (with a median of 680 M⊙ yr-1), efficient baryon to stellar mass conversion (20%), substantial gas reservoir and dust content, suggest rapid evolution and imminent quenching. Six of these galaxies reside in overdense filaments, while the remaining four may trace new distinct structures which will have to be spectroscopically confirmed. These few dusty galaxies dominate the star formation within the overdensity, contributing more than the numerous Hα emitters, and surpassing the cosmic average star formation rate density for this epoch. The presence of numerous very massive, dusty, and intensely star-forming galaxies at z5.2 shows that rapid stellar and dust mass assembly was already underway within the first billion years of cosmic history in overdense environments. Their properties, likely driven by efficient gas inflows along cosmic filaments in protocluster regions, suggest an accelerated evolution that current models and simulations have difficulty reproducing.
Key Words.:
Galaxies: high-redshift – Galaxies: evolution – Galaxies: starburst – Radio lines: galaxies – Methods: observational1 Introduction
At high redshift (z4), and despite recent progress, the contribution of dust-obscured galaxies to the young Universe remains quite uncertain (e.g., Traina et al. 2024). Recent ALMA surveys (as ALPINE or REBELS) showed that dust-obscured (far-IR, mm) and unobscured (UV rest-frame) star formation could have a similar contribution to the cosmic budget even at z6 (e.g., Gruppioni et al. 2020; Fudamoto et al. 2020; Khusanova et al. 2021; Barrufet et al. 2023; Fujimoto et al. 2024). In addition, recent observations by the JWST seem to show that heavily obscured sources do exist up to z12 (Rodighiero et al. 2023; Akins et al. 2023; McKinney et al. 2023; Zavala et al. 2023). These studies are not based on volume-complete surveys and/or rely on complex corrections and selections. The population of luminous and dusty galaxies at z4 is largely unconstrained and quantifying accurately the total density of the star formation rate (SFRD) at z4 from a well-controlled volume-limited survey stays an ambitious goal. One of the main contributor to the obscured SFRD at such high redshift are mm-bright, ultra massive, and optically-dark galaxies111Such galaxies are not detected in the deepest HST images (with typical 5-sigma depth of 27 mag for H band) but are detected with IRAC, and now routinely by the JWST.. A recent JWST study based on three mm-bright sources (Xiao et al. 2024) at z5-6 shows that they account for 20 % of the SFRD derived from rest-frame FUV to optical and corrected for dust attenuation at . These three galaxies are also ultra massive (log(M/M11.0) and would require, on average, about 50% of the baryons in their halos to be converted into stars. This is two to three times higher than even the most efficient galaxies at later times (Xiao et al. 2024, and references therein).
The two ultra-massive galaxies in the GOODS-N field (S3 and GN10) have been found to be located in a large-scale structure in the process of formation (Herard-Demanche et al. 2025), which also contains the well known HFD850.1 galaxy. HDF850.1 is forming in one of the richest environments identified to date at z5, with 100 z=5.17-5.20 galaxies (Herard-Demanche et al. 2025; Sun et al. 2024, see also previous studies e.g. Calvi et al. 2021). Observations suggest that such (massive) overdensities are not rare at (e.g. Guaita et al. 2022; Helton et al. 2024a). They are likely progenitors of today’s massive galaxy clusters. Understanding their distribution and properties shortly after the epoch of reionization helps bridge observations of the early Universe with modern-day structures. Given the existence of a similar galaxy overdensity in both GOODS-N and GOODS-S field, Sun et al. (2024) suggest that 50%20% of the cosmic star formation at z = 5.1-5.5 occur in protocluster environments.
Galaxies in high-redshift overdensities often have different properties compared to those in less dense regions. For example, Lemaux et al. (2022) found a highly significant positive relation between SFR and galaxy densities at , with galaxies in the densest proto-cluster environments exhibiting average SFR0.5 dex higher than those in the coeval field.
At higher redshifts (), Li et al. (2024a) observed that galaxies in dense environments exhibit a slightly elevated SFR at a given mass compared to those in less dense regions. Additionally, they found that galaxies in high-density regions tend to have redder UV slopes, which may indicate increased dust extinction (see also Helton et al. 2024b). In the GOODS-N z5.2 overdense region, Herard-Demanche et al. (2025) compare the properties of galaxies forming within the overdensity to those outside and find that the masses, star formation rates (SFRs), and UV luminosities of galaxies inside the overdensity are significantly higher than those outside.
Morishita et al. (2025) identified two overdensities of galaxies at z5.7 along the sightline of the galaxy cluster Abell 2744, where some members exhibit properties suggestive of accelerated maturation.
These studies show that the environment is influencing galaxy growth early in cosmic history even if high-z galaxies are also much more gas rich than local ones, for the same total mass, which could also give rise to very high SFR.
As part of the NIKA2 Cosmological Legacy Survey (N2CLS), we conducted very deep observations in GOODS-N at both 1.2 and 2 mm (Bing et al. 2023; Berta et al. 2025). We covered 160 arcmin2, to a depth close to the confusion noise of the IRAM 30 m telescope (Ponthieu et al. 2025). The large-scale overdense region at z=5.1-5.3, as observed by the FRESCO and JADES JWST programs (Herard-Demanche et al. 2025; Sun et al. 2024), is completely covered by N2CLS.
The survey gives a unique opportunity to study in detail the obscured star formation in the over-dense environment and analyse the properties of its bright-mm DSFG. The paper presents in particular the discovery of N2GN_1_13 at zspec=5.182, a massive DSFG belonging to the overdensity that we characterised using NOEMA observations. Furthermore, by combining our analysis with JWST/FRESCO data, we revealed another N2CLS DSFG at zspec=5.201. We additionally identify five N2CLS galaxies with a photometric redshift of z5.3, two of which are optically dark.
The paper is organized as follows. Section 2 gives a summary of the N2CLS survey in GOODS-N. In Sect. 3, we present the NOEMA observations of N2GN_1_13 and give its properties, that we then compare with HDF850.1 and GN10 in Sect. 4. In Sect. 5 (complemented by Appendices B and D), we present the remaining 7 N2CLS DSFGs identified at z (2 with zspec, 5 with zphot) and discuss their physical properties.
Section 6 is devoted to their efficiency for converting baryonic gas into stars. In Sect 7 we report their contribution to the cosmic SFRD (complemented by Appendix C) and present their localisation in the overdense structure.
In Sect 7, we compare our results with those expected from empirical, semi-analytical and hydrodynamical models. We finally provide a summary in Sect. 9. We also discuss in Appendix A the dust production in these high-redshift galaxies.
Throughout the paper, we adopt a CDM cosmology with km s-1 Mpc-1, and (Planck Collaboration et al. 2020), a Chabrier (2003) stellar initial mass function (IMF), and the Draine (2003) frequency-dependent dust absorption coefficient, , renormalized as indicated by Draine et al. (2014, i.e. m2 kg-1 as reference).
2 DSFG from the NIKA2 cosmological legacy survey in GOODS-N
N2CLS is the deepest and largest survey ever made at 1.2 and 2 mm, covering COSMOS and GOODS-N (Table 1 in Bing et al. 2023). In GOODS-N, it reaches an average 1 noise level of 0.17 and 0.048 mJy over 160 sq. arcmin. Our complete sample with a purity threshold 95% comprises 65 galaxies detected with SNR in a least one of the NIKA2 bands.
Thanks to a very successful NOEMA continuum observation program, combined with SMA 860 m (Cowie et al. 2017) and deep radio observations (Owen 2018), all GOODS-N sources have accurate positions that allow us to search for their multi-wavelength counterparts (Berta et al. 2025, hereafter B25). The final sample (including multiplicities) comprises 71 individual millimeter sources. Three galaxies (two identified with NOEMA, one identified with VLA) have no multi-wavelength counterpart and thus no redshift. We refer the reader to B25 for a complete description of the identification process and construction of multi-wavelength spectral energy distributions. We have spectroscopic redshifts for 29 sources222This includes the redshift measured with new NOEMA observations for N2GN_1_13 presented in this paper. (i.e. 41% of the sample) and photometric redshifts for 39 sources.
Among the 71 observed galaxies, 16 are classified as HST-dark (some of which are also IRAC-dark). Two galaxies remain unidentified. For the other 14 galaxies, the average redshift is z=4.9 and the average stellar mass is M⋆= 1.91011 M⊙ (stellar masses are derived using CIGALE, see Sect 3.2). This is in agreement with the general consensus in the literature that the majority of these optically-dark galaxies are heavily dust-obscured, star-forming sources at z3-6 that dominate the high-mass end of the galaxy population at these cosmic times (e.g., Alcalde Pampliega et al. 2019; Wang et al. 2019; Barrufet et al. 2023; Xiao et al. 2023; Gottumukkala et al. 2024).
3 N2GN_1_13: Identifying a new DSFG in the overdense environment of HDF850.1
We have selected one of these optically-dark N2CLS galaxies, N2GN_1_13, observed with JWST/NIRCam as part of the FRESCO program (Oesch et al. 2023) but without a confirmed redshift, as a secure target for NOEMA follow-up observations. N2GN_1_13 is a bright DSFG, with NIKA2 fluxes of 1.950.25 and 0.490.08 mJy at 1.2 and 2 mm, respectively. The accurate position of N2GN_1_13 has been secured thanks to our previous NOEMA program (W21CV, P.I. L. Bing), that detects at 2 mm a single counterpart with a flux of 39347 Jy, and a beam size of 1.61”0.93”. As illustrated in Fig. 1, the galaxy is absent in the HST images and only faintly visible in the IRAC images. The galaxy has a 6.2 single line detection in the NIRCam grism spectrum (Xiao, priv. comm.) but no conclusive redshift measurement. Assuming the NIRCam line is the Hα line, then z=5.196 and N2GN_1_13 would have the same redshift as the gigantic overdensity (Herard-Demanche et al. 2025; Sun et al. 2024).
Line | Fobs | FWHM | S | L | L |
[GHz] | [km ] | [Jy km ] | |||
CO(4-3) | 74.583 | 470110 | 0.270.08 | 1.650.49 | 5.191.54 |
CO(5-4) | 93.224 | 470⋆ | 0.370.10 | 1.450.39 | 8.892.40 |
⋆: The linewidth of the CO(4-3) line was adopted for the CO(5-4) line.
3.1 NOEMA observations
We observed N2GN_1_13 with NOEMA in band-1 using the C-configuration as part of program W23CX (PI: G.Lagache). The total on-source integration time was 6.6 hours. The receivers were tuned to cover a frequency range suitable for detecting the CO(4-3) and CO(5-4) transitions at z=5.196.
The NOEMA data were processed using the standard calibration and reduction procedures implemented in the GILDAS333https://www.iram.fr/IRAMFR/GILDAS/ software package. This resulted in calibrated uv-tables for the lower (LSB) and upper (USB) sidebands, sampled with a spectral resolution of 20 MHz. The source 1125+596 was used as the phase and amplitude calibrator. The absolute flux uncertainty is estimated to be within 10%, while the positional accuracy is better than 0.16′′. The final dataset achieves a synthesized beam of , with a line sensitivity of 0.7 mJy/MHz and a continuum sensitivity of 5.7 Jy/beam.
We detect two spectral lines with SNR ¿ 6, which we unambiguously identify as CO(4-3) and CO(5-4) at z=5.182. As expected from the CO/[CI] flux line ratio in distant galaxies and our sensitivity, the [CI](1-0) at 79.6 GHz remains undetected with S Jy km (assuming the same FWHM as CO(4-3))444The 3 upper limit was computed as , where and are the RMS noise and the velocity width of a spectral channel, respectively.. Figure 1 presents the two CO detections, fitted along with the continuum. The best-fit Gaussian parameters derived from the line profiles are listed in Table 1.
To characterize the continuum emission, we combined all channels from the LSB and USB subbands, excluding the spectral windows containing line emission, and produced a double sideband (DSB) continuum map (Fig. 1). The map was cleaned using natural weighting with a limited number of clean components. The continuum flux was measured using two methods: integration within the 2 contour and direct measurement at the position of the emission peak. Both approaches yield a flux of 576 Jy at 84.6 GHz, suggesting that the continuum source is unresolved.
3.2 Properties of N2GN_1_13
The total IR luminosity (integrated between 8 and 1000 m), SFR, stellar mass , dust mass and dust spectral index have been derived in B25 using multi-wavelength SED fitting.
The dust mass has been obtained using Draine (2003) frequency-dependent dust absorption coefficient, , renormalized as indicated by Draine et al. (2014), i.e. m2 kg-1 as reference. The panchromatic optical to radio SEDs are modeled with the MAGPHYS and SED3FIT codes, in their original and high- versions (da Cunha et al. 2008, 2015; Battisti et al. 2020; Berta et al. 2013), as well as with CIGALE (Burgarella et al. 2005; Noll et al. 2009; Boquien et al. 2019). The spectral index is obtained using a modified black-body fit to the far-IR to mm data. The effect of the cosmic microwave background is taken into account as described by da Cunha et al. (2013).
The stellar mass has been obtained using the same assumptions for CIGALE and the high- versions of MAGPHYS and SED3FIT, i.e. Bruzual & Charlot (2003, BC03) stellar populations with a Chabrier (2003) IMF and the Charlot & Fall (2000) two-components attenuation law. B25 presents the comparison of stellar masses obtained using the different codes, and shows a very good agreement, but for few outliers lying at high z. Unfortunately, the sources considered in this paper are among the outliers. For example, considering the 7 optically-dark sources at z5.2 (see Table 3), the ratio between the high- versions of MAGPHYS or SED3FIT and CIGALE stellar mass determination ranges from 1.4 to 9.2, with an average ratio of 5.2.
Masses up to 1012 M⊙ are derived using MAGPHYS and SED3FIT (and CIGALE for N2GN_1_20), indicating the presence of extremely massive objects at this redshift.
Excluding any AGN contamination, the high flux observed in the long-wavelength NIRCam/JWST F444W filter (0.7-1.3 Jy) supports by itself the very large stellar masses, i.e. M1011-1012 M⊙. Such very high stellar masses goes with very large extinction.
For N2GN_1_13, the stellar masses from SED3FIT and CIGALE are log(M⋆)=11.68 and log(M⋆)=11.03, respectively. In the following, and to be conservative, we adopt the lower stellar mass value obtained using CIGALE. To be consistent, all the other parameters (e.g. SFR, dust mass) are taken from the CIGALE SED fitting.
The total molecular gas mass is obtained using
Following Berta et al. (2023) (see also Dunne et al. 2022), we adopt a conversion factor = 4.0 M⊙ (K km s-1 pc2)-1, including the contribution of helium. The luminosity of the CO lines detected by NOEMA are transformed into
adopting the average ratios given by Carilli & Walter (2013) for SMGs, , and , which give (and also , see Sect. 4) consistent with our measurement. We obtain = (2.30.6)1010 M⊙ and (3.10.9)1010 M⊙ from the CO(5-4) and CO(4-3) line luminosity, respectively. We adopt the molecular gas mass from the CO(4-3) which is detected with a higher SNR and with a measurable FWHM (Table 1).
We can now derive the depletion time of the molecular gas in each galaxy using the molecular gas mass and the SFR:
This is the time it would take for a galaxy to consume the entirety of its molecular gas reservoir, if it formed stars at the rate derived from far-IR and sub-mm data in a single event (not considering any mass return to the ISM from stellar winds and supernovae).
We also finally compute the gas-to-dust mass ratio GDR=MMol/Mdust.
We summarise the physical properties of N2GN_1_13 in Table 2 and show JWST/NIRCam images in Fig 7. Based on these findings, we can conclude that:
N2GN_1_13 has a very high SFR and is thus in the process of very efficient stellar mass build-up. It has a specific star formation rate log(sSFR)=-8.18. Compared to the main sequence of star formation, N2GN_1_13 lies above, with SFR = log(SFR/SFRMS)= 0.55 (if we use the parametrization of the MS from Popesso et al. 2023) and SFR = 0.29 (if we use the parametrization of Speagle et al. 2014). The former exhibits the well-known flattening at large stellar masses, whereas the latter does not.
N2GN_1_13 has a very red continuum with ABmag. As shown in B25, the optical photometry reveals a rest-frame UV excess, which is reproduced by adding an unextinguished young (ages between 10 and 100 Myr) simple stellar population to the SED modeling. As can be seen on Fig. 7, it has an extended morphology, in contrast with the so-called little red dots, LRD (e.g., Labbé et al. 2023; Greene et al. 2024; Matthee et al. 2024; Pérez-González et al. 2024). The source is very red simply because it is dusty with i) a very high Av=3.960.13 mag, indicating a large dust content and ii) rest-frame dust emission detected by NOEMA at both 324 and 574 m. Furthermore, this source cannot be classified as a broad-line AGN, as the line does not exhibit broadening. It has no X-ray counterpart and no rest-frame mid-IR data are available to verify the possible presence of an AGN-torus component contributing to its mid-IR SED.
The very high SFR relative to the available gas reservoir results in a short depletion time of 43 Myr, indicative of a brief (100 Myr) starburst phase555Adopting = 1 M⊙ (K km s-1 pc2)-1 (e.g., Riechers et al. 2020) would give a depletion time 4 times lower, making this source even more extreme. The 3 upper limit on the CI(1-0) line gives M1010M⊙ and 120 Myr, assuming a conversion factor = 17 M⊙ (K km s-1 pc2)-1 from Dunne et al. (2022).. Such a low points to an extreme case of star formation. Assuming no gas accretion or replenishment, and modeling the gas consumption as an exponential decay , the galaxy’s SFR would drop to 10% of its initial value within approximately 100 Myr, effectively marking the quenching of star formation.
With a dust mass of 7.70.8108 M⊙, and given its stellar mass, N2GN_1_13 lies near the maximal theoretical dust yield expected from the combined contributions of supernovae (assuming no reverse-shock destruction) and AGB stars (see Fig. 5 and Witstok et al. 2023). However, the short gas depletion time suggests that dust production from AGB stars, whose contribution operates on longer timescales, is likely minimal. This implies that additional dust formation mechanisms, such as rapid grain growth in the ISM, may be required to account for the observed dust mass.
Our observed gas-to-dust ratio leads to a galaxy metallicity consistent with solar, (e.g., Li et al. 2019). However, this estimate should be interpreted with caution, as our adopted values for and are themselves based on the assumption of solar-like metallicity.
With a CO(5–4)/CO(4–3) flux ratio of 1.37, N2GN_1_13 closely matches the ratio observed in the galaxy CGG-z4.b at (Brinch et al. 2025), and falls within the upper range of values reported in the literature (see the Fig. 5 in Brinch et al. 2025). This elevated ratio may suggest the presence of high gas temperatures, densities, and/or pressures.
N2GN_1_13 is thus a massive, extended, and highly obscured galaxy caught in a short-lived yet extreme starburst phase at z5.2. Its high stellar mass, red optical continuum, and gas reservoir point to rapid evolution and imminent quenching.
Quantity | Unit | N2GN_1_13 | N2GN_1_06 | N2GN_1_01 |
HDF850.1 | GN10 | |||
z=5.182 | z= 5.183 | z=5.303 | ||
LIR | 1012 L⊙ | 6.50.4 | (8.20.6) | 15.00.1 |
SFR | M⊙ yr-1 | 70341 | (89968) | 163982 |
1.900.21 | 2.560.36 | 2.980.06 | ||
Mdust | 108 M⊙ | 7.70.8 | (11.91.5) | 16.70.8 |
M⋆ | 1011 M⊙ | 1.10.1 | (0.60.2) | 1.60.8 |
MMol | 1010 M⊙ | 3.10.9 | (125) | 253 |
= MMol/SFR | Myr | 4313 | 13754 | 15520 |
GDR= MMol/Mdust | 4012 | 10342 | 15220 |
4 Gas masses, gas-to-dust ratios, gas depletion times for the 3 brightest DSFGs in the overdense environment
In addition to N2GN_1_13, two other galaxies in the overdense environment have CO measurements: GN10 and HDF850.1 (see the compilation in Riechers et al. 2020, R20). We use the same recipes as in Sect. 3 to compute gas masses, GDR and gas depletion times for these two galaxies and use the other source properties (e.g., SFR) as derived from the SED fitting with CIGALE. The results are given in Table 2.
We use the CO(2-1) line measurements for HDF850.1 and GN10 and the conversion factor from Carilli & Walter (2013), which gives MMol consistent with R20 if we would consider =1 rather than =4, and assuming a mean gravitational lensing magnification factor =1.6 for HDF850.1 as in R20. If we further correct for (which is equal to 0.057 m2 kg-1 in R20 versus 0.047 here), we would obtain GDR=46 and 31 w.r.t. 65 and 130 in R20, and =39 and 34 Myr w.r.t to 70 and 40 Myr in R20, for GN10 and HDF850.1, respectively.
The large difference in GDR for HDF850.1 arises from a factor 3.7 difference in Mdust (assuming the same ), while the discrepancy in for GN10 is mainly due to a factor 1.6 in SFR. Notice that we have the same spectral index as in previous studies (R20, Walter et al. 2012).
We derive gas mass fraction fgas = MMol / (MMol+M⋆) of 0.6 for HDF850.1 and GN10, which is typical of gas fraction of z5 sub-millimeter galaxies (Dudzevičiūtė
et al. 2020). For N2GN_1_13, the gas mass fraction is 3 times lower but remains consistent with certain z5 SMGs reported in the literature (e.g. Zavala et al. 2022). However, we warn that such measurements have to be taken with caution.
For instance, Sun et al. (2024) reported a gas fraction of 0.18 for HDF850.1, which is significantly lower than the value obtained in this study. This discrepancy primarily arises from the considerably lower gas mass reported in Neri et al. (2014), attributed to the adoption of a much lower (this is also the case in Zavala et al. (2022) where =1.0 M⊙ (K km s-1 pc2)-1).
Tacconi et al. (2020) give updated scaling relations that links M⋆, redshift and distance from the main sequence of star forming galaxies (from Speagle et al. 2014) to molecular gas depletion timescale () and thus molecular gas mass (M). We use these relations to compute and M (see Table 4) and obtain = 206, 142, 161 Myr and M=1.41011, 4.71010, 2.41011 M⊙ for N2GN_1_13, HDF850.1 (delensed) and GN10, respectively. While this compares quite well to the results given in Table 2 for HDF850.1 and GN10 (10% level), M (and thus ) is overestimated by a factor 4.3 for N2GN_1_13. If we were using higher J transitions to get the molecular mass for HDF850.1 and GN10, we would have obtained the same discrepancy (i.e. molecular gas mass of 3.1 and 5.41010M⊙ using Jup=5 rather than 12 and 251010 M⊙). This shows how these determinations have to be taken with caution.
5 Bright dusty galaxies in the overdense environment
NIKA2 ID | Alt.† | RA | Dec | z | z | Comment |
[hh:mm:ss] | [dd:mm:ss] | |||||
N2GN_1_01 | GN10, S2 | 12:36:33.44 | 62:14:08.7 | 5.303s | Optically-dark | |
N2GN_1_06 | HDF850.1 | 12:36:51.99 | 62:12:25.7 | 5.185s | Optically-dark | |
N2GN_1_13 | GN32 | 12:36:58.61 | 62:14:51.1 | 5.182s | Optically-dark | |
N2GN_1_23 | GN15, S3 | 12:36:56.54 | 62:12:07.5 | 5.179s | Optically-dark | |
N2GN_1_61 | P-40 | 12:36:57.48 | 62:16:54.4 | 5.201s | Optically-dark | |
N2GN_1_20 | AzGN4 | 12:35:50.22 | 62:10:42.4 | 5.330.32p | 5.16 | Optically-dark |
N2GN_1_43 | 12:37:11.95 | 62:11:19.7 | 5.870.38p | 5.31 | Optically-dark | |
N2GN_1_26 | GN38 | 12:37:41.14 | 62:12:20.4 | 5.310.08p | 4.87 | |
N2GN_1_33 | GN25 | 12:36:29.03 | 62:10:45.5 | 5.390.07p | 5.29 | |
N2GN_1_57 | HRG14 | 12:37:31.69 | 62:16:16.5 | 5.300.18p | 5.18 |
†: The HDF850, GN, AzGN, HRG14, S and P- nomenclatures were defined by Hughes et al. (1998), Pope et al. (2005), Perera et al. (2008), (Finkelstein et al. 2015), Xiao et al. (2024), Greve et al. (2008).
‡ s: spectroscopic redshift; p: photometric redshift from B25.
h photometric redshift from New-Hyperz
5.1 N2CLS galaxies at similar redshifts
In addition to N2GN_1_01 (GN10), N2GN_1_06 (HDF850.1) and N2GN_1_13 discussed above, we have 7 more galaxies with redshifts compatible with the overdense structure (see Table 3).
Two are optically-dark N2CLS galaxies with spectroscopic redshift from JWST: N2GN_1_23 (also known as S3 in (Xiao et al. 2024) and N2GN_1_61.
For the five additional galaxies with previously reported photometric redshifts of (see B25 and references therein), independent photometric redshifts were also derived using the enhanced JWST photometric coverage (when available) and the SED-fitting code New-Hyperz. This is a modified version of the public code Hyperz originally developed by Bolzonella et al. (2000)666https://people.osupytheas.fr/~rpello/newhyperz/. The procedure and its validation is discussed in Appendix B. Results are shown in Table 3. Our new photometric redshifts are within one-sigma uncertainty of the literature based , for all but N2GN_1_26 (2 ). The new photometric redshifts support their association with the overdensity.
Among these five N2CLS galaxies, two are optically dark (N2GN_1_20 and N1_1_43), while the remaining three are optically bright (N2GN_1_26, N2GN_1_33, N2GN_1_57).
Details on each source are provided in Appendix D.
On the 10 sources (5 with and 5 with ), N2GN_1_01, 06 and 23 are already identified to belong to the overdensity in Sun et al. (2024); Herard-Demanche et al. (2025); Xiao et al. (2024) ; N2GN_1_13 and N2GN_1_61 fall into their area of study but were not previously identified. ; N2GN_1_33 and 43 are in the FRESCO and JADES footprint, respectively, but were not reported ; N2GN_1_57 lies at the border of the JADES footprint and N2GN_1_20, 26 are further away.
NIKA2 ID | SFR | AV | M⋆ | Mdust | Mgas | |
[M⊙ yr-1] | [1010 M⊙] | [108 M⊙] | [1010 M⊙] | [Myr] | ||
N2GN_1_01 | 163982 | 4.00.0 | 15.80.8 | 16.70.8 | 24.3 | 160.8 |
N2GN_1_06 | 36027 | 3.70.2 | 2.30.8 | 4.70.6 | 4.7 | 141.9 |
N2GN_1_13 | 70341 | 4.00.1 | 10.71.2 | 7.70.8 | 13.7 | 206.2 |
N2GN_1_23 | 51675 | 3.10.1 | 7.60.6 | 5.90.9 | 11.1 | 190.3 |
N2GN_1_61 | 18460 | 2.1 0.5 | 1.20.8 | 4.53.5 | 3.2 | 134.3 |
N2GN_1_20 | 2160304 | 2.70.2 | 30340 | 23.313.4 | 102.9 | 476.3 |
N2GN_1_43 | 658112 | 3.40.2 | 10.71.9 | 6.50.7 | 12.4 | 197.6 |
N2GN_1_26 | 196298 | 1.30.0 | 9.30.5 | 18.30.9 | 20.3 | 118.0 |
N2GN_1_33 | 1874101 | 1.70.0 | 10014 | 5.50.9 | 59.2 | 316.1 |
N2GN_1_57 | 36918 | 2.20.0 | 6.00.3 | 4.20.7 | 7.2 | 231.5 |
† corrected from =2.5. |
5.2 Source properties
The multi-wavelength SEDs for these 10 sources are shown in B25. Using CIGALE (see Sect. 3.2), we derive their properties that are summarised in Table 4. Our sample is characterised by high V-band attenuation (AV), SFR and stellar mass, with median values of AV=2.93, SFR=681 M⊙ yr-1 and log=11. These values are comparable to those reported for the three ultra-massive galaxies in Xiao et al. (2024)
However, they are systematically much higher than those found in the larger DSFG sample from JWST (36 galaxies) analyzed in the same study.
Combining the derived SFR and , and the known redshift, we apply the Tacconi et al. (2020) scaling relations to derive the molecular gas depletion time scale, , and mass, of these N2GN sources. Figure 2 locates the sources in the vs. SFR plane and compares their position to the main sequence (MS) loci defined by Speagle et al. (2014) and Popesso et al. (2023). The Tacconi et al. (2020) scalings are based on the distance of the sources from the Speagle et al. (2014) MS, which does not include the bending at large , appearing clearly instead in Popesso et al. (2023).
The N2GN z5.2 galaxies lie significantly above the star-forming main sequence (MS) as defined by Popesso et al. (2023), indicating a strongly bursty mode of star formation. Although they appear closer to the MS when adopting the Speagle et al. (2014) parametrization, eight galaxies still lie above the 1 envelope of the MS.
The median obtained for these galaxies is of 0.19 Gyr, with individual values spanning over the range between and Gyr. These short time scales (or equivalently very high star formation efficiency, SFE, yr-1) confirm the intense ongoing starbursts activity that powers their FIR and mm emission. Despite their short depletion timescales, the starbursts are still operating over periods significantly longer than their dynamical timescales, suggesting that the activity may be sustained through multiple bursts or extended starburst phases. The fact that these galaxies reside in a confirmed overdense environment likely provides both the fuel and favorable conditions to maintain such prolonged star formation.
Given their short gas depletion timescales, these DSFGs may transition into a quiescent phase by , assuming a constant SFE, complete molecular gas consumption, and no external gas replenishment. This evolutionary pathway would be consistent with the emergence of massive quiescent galaxies observed in other overdense environments at (e.g., Carnall et al. 2024; Kakimoto et al. 2024; Tanaka et al. 2024).
Figure 2 also displays the locations of the DSFGs identified by Xiao et al. (2024), selected based on their JWST colors mag, which suggest significant dust attenuation. Notably, these JWST-selected galaxies lie systematically below the star-forming main sequence. This offset may stem from underestimated SFRs and/or overestimated M⋆. In the absence of far-infrared or (sub-)millimeter constraints (”No FIR” case), UV-based SFRs, corrected for dust attenuation, frequently fail to capture the bulk of star formation in dusty systems. GOODS-ALMA studies (Elbaz et al. 2018; Xiao et al. 2023) report typical ratios of total to dust-corrected UV SFRs on the order of 8. Conversely, Pacifici et al. (2023) find that including FIR data in SED fitting yield systematically lower SFRs compared to fits that omit such data. To investigate this further, we performed SED fitting on a subset of our N2CLS DSFGs using only optical-to-NIR photometry (m). Using BAGPIPES (Carnall et al. 2018), we modeled the UV-to-NIR SEDs of six sources with available HST+JWST photometry (N2GN_1_01, 06, 13, 23, 33, and 43). For N2GN_1_01 (S2) and N2GN_1_23 (S3), we confirm the results of Xiao et al. (2024), with SFRs significantly underestimated. However, this is not the case for N2GN_1_33 and N2GN_1_43, where higher SFRs are recovered. To further probe this discrepancy, we repeated the analysis using the CIGALE code. The inferred SFRs from the No IR fits differ notably from those obtained with the full UV-to-(sub)mm coverage, with differences reaching up to a factor of 30. SFRs are overestimated in the No IR scenario, except for two sources. Importantly, the stellar masses derived in the No IR case are systematically overestimated, with offsets ranging from a factor of 1.1 up to 7.5. These comparisons underscore that, even with the unprecedented depth of JWST, (sub-)mm data remain essential to accurately determine SFRs, M⋆ (and AV) in dusty galaxies. However, such data require very deep observations. Indeed, none of the 25 DSFGs in the GOODS-N field reported by Xiao et al. (2024) (except for the brightest sources N2GN_1_01, 06, and 23) are detected at in N2CLS, which already reaches the confusion limit in GOODS-N (Ponthieu et al. 2025).
We finally examined the rest-frame optical morphologies of our sources using JWST/NIRCam imaging (Fig. 7). Most galaxies appear extended or clumpy, with no strong evidence for point-like or compact morphologies—only one galaxy in the sample appears marginally compact. Several galaxies show irregular or asymmetric light distributions, suggestive of ongoing or recent dynamical activity, such as minor mergers, interactions, or internal disk instabilities. These features are broadly consistent with recent JWST results for high-redshift, dust-obscured galaxies (e.g., McKinney et al. 2025), which reveal a wide range of morphologies, from clumpy and irregular to disk-like systems, among massive DSFGs.
6 Stellar baryon budget and the efficiency of star formation
Recent observations have unveiled a substantial population of massive galaxies in the early universe, challenging existing models of galaxy formation. Studies leveraging data from the JWST have identified galaxies at redshifts to with stellar masses approaching 1011 M⊙. For instance, utilizing JWST/MIRI data, Wang et al. (2024) found high densities of massive galaxies at , implying elevated stellar baryon efficiencies . Specifically, they report that the most massive galaxies at 8 require , compared to , for typical low-redshift galaxies.
Similarly, Xiao et al. (2024) examined 36 massive, dust-obscured galaxies with spectroscopic redshifts between and from the JWST FRESCO survey. They found that three ultra-massive galaxies (log M⋆/M11.0, among them S2N2GN_1_01 and S3N2GN_1_23) require an exceptional of approximately 50%, which is two to three times higher than the efficiencies observed in galaxies at later epochs.
Further, Gentile et al. (2024) reported on two massive and dusty starburst galaxies at z discovered in the COSMOS-Web survey. Their analysis indicates of approximately 0.25 and 0.5, respectively.
In numerical simulations, is found to increase significantly with both stellar mass and redshift, indicating that, at fixed mass, halos at (very) high redshifts convert gas into stars more efficiently than their lower-redshift counterparts (e.g., Ceverino et al. 2024).
All these results imply that certain populations of high-redshift galaxies may experience very efficient star formation at early times.
Building upon these studies, we have computed the stellar baryon conversion efficiency for our sample of galaxies at , following the methodology adopted in Boylan-Kolchin (2023); Xiao et al. (2024).
Using the cumulative halo mass function, we compute the comoving number density of halos above a given halo mass threshold, at redshift . The maximum stellar mass () that is statistically allowed within an observed volume is given by , where represents the (integrated) efficiency with which baryons are converted into stars and =0.156 is the cosmic baryon fraction from Planck Collaboration et al. (2020). We show on Fig. 3 our 10 galaxies in the stellar mass vs. redshift plane. The lines correspond to different number densities:
3.710-6 Mpc-3 (i.e. one source in the cosmic volume explored by N2CLS, assuming z=5.20.5 ; with and displayed as blue-dotted and grey-dashed lines, respectively) and 3.910-12 Mpc-3 (i.e. a single galaxy in the whole sky; with displayed as a continuous grey line).
Our analysis reveals notably high values (0.2 for all but N2GN_1_06 and 61), indicating that a significant fraction of baryonic mass in these galaxies has been converted into stars.
Two galaxies (NN2GN_1_20 and 33), identified through photometric redshifts, lie above the model expectations for the GOODS-N survey volume. This discrepancy could arise from an overestimation of their photometric redshifts, assumptions related to the IMF (for example, a top-heavy IMF could yield lower stellar masses by dex, resulting in slightly reduced values of ) or from the presence of an unusually dense concentration of massive halos in this region of the sky.
The elevated values of observed in our sample suggest that these galaxies are undergoing unusually efficient star formation at . Given their probable high metallicities and host halo masses exceeding , it is unlikely that cold gas inflows are the primary driver of this efficiency. Instead, the observed efficiency may reflect a temporary suppression of stellar feedback, as proposed in feedback-free starburst scenarios for higher-redshift galaxies with lower metallicities (e.g., Dekel et al. 2023; Li et al. 2024b), or merger-driven inflows that channel gas into central regions and ignite intense starbursts, or compaction events that funnel enriched gas into central regions, triggering intense, short-lived starbursts.
7 Spatial distribution and abundance of N2CLS DSFGs in the GOODS-N overdensity
7.1 Contribution to the star formation rate density at
Recent studies have shown that a significant fraction of the cosmic star formation rate density (SFRD) at high redshift can be driven by rare, massive galaxies with high star formation efficiencies. For instance, Xiao et al. (2024) found that at , ultra-massive galaxies with can account for up to % of the total SFRD (from Madau & Dickinson 2014). These findings highlight the critical role of massive, DSFGs in early cosmic star formation, motivating a closer look at their contribution in overdense environments such as GOODS-N.
The combined cosmic SFRD from the five N2CLS galaxies with spectroscopic redshifts777Assuming an area of 160 arcmin2 and to compute the volume. is M⊙ yr-1 Mpc-3, representing a factor of 1.390.06 excess relative to the total SFRD reported by Madau & Dickinson (2014) at similar redshift. Including the additional contribution from galaxies with photometric redshifts increases the total SFRD to M⊙ yr-1 Mpc-3, corresponding to a factor of 4.270.16 higher than the cosmic average.
The values are shown in a figure in Appendix 6.
The elevated SFRD reflects the fact that our sample is embedded within a confirmed overdense environment, where star formation is significantly enhanced relative to typical field regions at this epoch.
More broadly, current research suggests that massive galaxy overdensities at high redshift may be relatively common (e.g., Kashino et al. 2023; Wang et al. 2023; Helton et al. 2024a; Brinch et al. 2024). In the GOODS-N field, Lin et al. (2025) estimate the fraction of the cosmic SFRD hosted within the protocluster environment at z5.2, based on the number density of H emitters. They find that approximately 55% of the SFRD at this redshift is associated with the protocluster, somewhat higher than the 30% predicted by simulations (Chiang et al. 2017) but in good agreement with recent results based on the FLAMINGO simulation suite (Lim et al. 2024). Within the same volume considered in Lin et al. (2025) (i.e., 62 arcmin2 and ), our four DSFGs yield a star formation rate density of [M⊙ yr-1 Mpc-3]) = . This value is 1.6 times higher than the estimate based on the 150 H emitters alone, suggesting that a substantial fraction of star formation activity is missed even in deep JWST surveys of highly overdense environments.
7.2 Environmental localization of N2CLS DSFGs in the GOODS-N overdensity
Recent JWST observations from the JADES and FRESCO programs have revealed that the overdensity surrounding HDF850.1 represents one of the richest environments identified to date at , with approximately 100 galaxies at =5.17–5.20 distributed across 10 substructures within a (15 cMpc)3 volume (Herard-Demanche et al. 2025; Sun et al. 2024). This region is particularly extreme, as 28% of all –6.0 H emitters detected across the GOODS-N field are concentrated within a narrow redshift slice of . The spatial extent of this protocluster is consistent with the typical size of such structures predicted by Chiang et al. (2017) using a set of N-body simulations and semi-analytic models. In addition, Calvi et al. (2023) identified the galaxy GN10 (N2GN_1_01) as residing in an overdense region based on photometric redshifts, a result that has now been confirmed spectroscopically by Sun et al. (2024).
We show the locations of the N2CLS galaxies within the overdensity in Fig. 4. Cosmological simulations suggest that at , approximately 90% of baryonic and dark matter resides in unvirialized, filamentary large-scale structures (e.g., Haider et al. 2016). Sun et al. (2024) interpret the 3D spatial distribution of group member galaxies as being consistent with three distinct filamentary structures. For simplicity, we compare the positions of our galaxies relative to the filaments identified in Sun et al. (2024), rather than to the 18 substructures detailed in Herard-Demanche et al. (2025). The –5.20 overdensity, traced by filament-1 and filament-2, spans the full extent of the GOODS-North FRESCO field (62 arcmin2). Within this region lie four N2CLS DSFGs with spectroscopic redshifts: N2GN_1_06 and N2GN_1_23 are located in filament-1, while the newly identified N2GN_1_13 and N2GN_1_61 fall within filament-2. In filament-3, N2GN_1_26 (photometric redshift) is found along with GN10, and N2GN_1_43 lies nearby. N2GN_1_57 may trace a distinct, possibly new, structure. We note that some previously confirmed spectroscopic members of the protocluster (Walter et al. 2012; Calvi et al. 2021) fall outside the JADES and FRESCO joint footprint, implying that the overdensity may exhibit more complex substructures than currently captured. The sources N2GN_1_20 and N2GN_1_33 are located farther from the core overdensity. Interestingly, Jiang et al. (2018) report a giant protocluster at occupying a volume of cMpc3, embedded within a larger overdense region spanning cMpc2 and extending 34 cMpc along the line of sight. Further investigation is needed to determine whether the overdensity extends toward the regions traced by N2GN_1_20 and N2GN_1_33.
The spatial distribution of N2CLS galaxies at z5.2, when viewed in two dimensions (RA, DEC), does not exhibit a recognizable filamentary pattern, unlike the tentative identification of a filament with NIKA2 in the GJ526 field (Lestrade et al. 2022).
8 Accelerated galaxy evolution in the overdense structure
Chiang et al. (2017) outlined three evolutionary stages of cluster formation: an early inside-out growth phase from to , an extended period of vigorous star formation from to , and a final stage of rapid mass assembly and quenching from to . Our results are consistent with the second phase of this sequence, during which intense star formation is expected to occur across 10–20 cMpc-scale structures, contributing significantly to the stellar mass of present-day clusters. Interestingly, Lim et al. (2024), using the FLAMINGO simulation suite, suggest that the suppression of star formation in protocluster environments may already begin by , indicating that this epoch could represent a transitional phase in cluster evolution.
Recent results from the FOREVER22 project (Morokuma-Matsui et al. 2025) indicate that the elevated star formation activity observed in protocluster galaxies at z6, compared to field galaxies, is primarily driven by a higher abundance of massive galaxies rather than enhanced specific star formation rates. This trend is likely a consequence of the denser environments in protoclusters, where increased galaxy interactions and sustained gas accretion promote the rapid assembly of massive systems. These massive galaxies tend to exhibit stronger dust attenuation, resulting in redder UV continuum slopes. This is consistent with the findings of Helton et al. (2024b), who reported that the UV-brightest and UV-reddest galaxies are preferentially located in denser environments with a higher number of neighboring galaxies. Together, these observations provide compelling evidence that galaxy evolution is accelerated within high-redshift overdense structures and may help explain the record number of bright DSFGs we identified in this particular structure.
In the FoREVER22 project, protocluster galaxies show a combination of stellar masses and star formation rates that are somewhat lower, but still comparable to those of the DSFGs identified in our study. This places our DSFGs at the upper end of the FoREVER22 population in terms of total star formation activity. In contrast, the FLAMINGO simulations from Lim et al. (2024) produce galaxies with high star formation rates that are consistent with the SFRs measured for our sample. Furthermore, Chiang et al. (2017) estimate that the total SFR in protoclusters at is typically below 1000 M⊙ yr-1 which is about four times lower than the combined SFR of our five spectroscopically confirmed DSFGs. This unusually high level of star formation highlights the exceptional nature of this structure.
In the context of the SIDES simulation (Béthermin et al. 2017; Gkogkou et al. 2023), where no environmental effects are considered, we can use this dataset to address whether the enhanced dusty star formation observed here is driven by clustering or by the presence of large gas masses. To explore this, we selected SIDES galaxies in the 2-deg2 simulation in the redshift range and with a NIKA2 1.2 mm flux greater than 0.7 mJy, which mimics the selection criteria for our sample (B25). The SFR–M⋆ relation for a NIKA2-like population is shown in Fig. 2. With the exception of N2GN_1_61, SIDES galaxies generally exhibit lower SFR and M⋆.
In SIDES, the SFR is capped at a hard limit of 1000 M⊙/yr, and each galaxy’s SFR is redrawn until it falls below this threshold.
This is explaining the lack of SFR1000 M⊙/yr galaxies in SIDES.
However, we can also notice that SIDES does not reach the high stellar masses. The most massive galaxy in SIDES has a stellar mass of log(M⋆) = 10.92. Even when considering the larger 117-deg2 SIDES-UCHUU simulation (Gkogkou et al. 2023), the most massive galaxy reaches log(M⋆) = 11.29.
SIDES employs abundance matching to obtain the stellar mass, which connects halo to galaxy stellar mass via mass functions. This approach may hinder the existence of high-stellar mass objects if they are not represented in the observed stellar mass function used to build the SIDES simulation.
Indeed, using the stellar mass function from SIDES, we expect 0.4 galaxies with log(M⋆/M10.7 at in the GOODS-N area (160 arcmin2), but we observe nine.
Given the non-uniform distribution of sources, we investigated the impact of field-to-field variance by dividing the SIDES-UCHUU simulation into 2,562 independent sub-fields, each covering 160 arcmin2. Among these, 13 sub-fields contain 4 galaxies above the stellar mass threshold, a small fraction but consistent with our spectroscopic () sample.
To evaluate the effect of stellar mass uncertainties, we added a random scatter of 0.5 dex to the SIDES stellar masses and repeated the sub-field analysis. With this added uncertainty, the average stellar mass of the most massive galaxy in the redshift range across all sub-fields is , consistent with the most massive spectroscopically confirmed galaxy in our sample.
Furthermore, we now identify 25 sub-fields (1% of the total) that contain 8 galaxies with within . In conclusion, while such overdensities are rare, their existence is plausible when considering both the strong clustering of massive galaxies at and the significant uncertainties in stellar mass estimates.
Based on their stellar masses, the five galaxies are each expected to reside in massive dark matter halos, with a combined halo mass of M⊙, as inferred from the empirical – relation at (Behroozi et al. 2019). While it is uncertain whether all these galaxies will eventually merge into a single structure by , such an outcome would imply a descendant halo more massive than that of a typical Coma-like cluster progenitor at the same redshift, which simulations suggest is closer to M⊙ (Chiang et al. 2013). This highlights the exceptional mass scale of the environment traced by these galaxies. However, it is important to note that total mass estimates for protoclusters are sensitive to the choice of aperture and redshift interval used for their definition, influencing such comparisons across studies (Lim et al. 2024).
9 Conclusion
Using 1.2 mm observations with NIKA2 in GOODS-N, we identify ten DSFGs at , many of which are extremely faint or even undetected in deep HST and some JWST imaging. These galaxies exhibit remarkably high SFRs, substantial dust content, and significant stellar masses, placing them among the most extreme star-forming systems known at this epoch. Six are securely associated with the exceptional overdensity reported in Sun et al. (2024) and Herard-Demanche et al. (2025).
We first identify a new DSFG in the overdense environment of HDF850.1 through NOEMA observations, which detect CO(4–3) and CO(5–4) lines at . In addition, another N2CLS galaxy shows a 7 emission line in FRESCO data, consistent with H at z. Combined with previously known galaxies (HDF850.1, GN10, and S3), this brings the number of spectroscopically confirmed DSFGs within the overdense structure to five. We further identify five DSFGs with photometric redshifts consistent with the overdensity.
These galaxies exhibit exceptionally high star formation rates (ranging from 200 to 2000 M⊙,yr-1) and large stellar masses (log(M⋆/M⊙) = 10–12), placing most of them significantly above the star-forming main sequence at , especially under the Popesso et al. (2023) parametrization. Their properties are comparable to those of the ultra-massive galaxies analyzed by Xiao et al. (2024), suggesting an early and efficient buildup of both stellar and dust mass. They show short gas depletion timescales (with a median of 0.19 Gyr) and remarkably high stellar baryon conversion efficiencies (20%, compared to 10% locally). The observed dust masses further support the scenario of efficient dust formation. Due to their short gas depletion timescales, these DSFGs could evolve into quiescent galaxies by , assuming no new gas inflow. This scenario aligns with the early formation of massive quiescent galaxies seen in similar dense environments.
Their contribution to the cosmic SFRD is significant, exceeding the total SFRD by a factor of 1.4 when considering spectroscopically confirmed galaxies, and by a factor of 4.3 when including all ten sources. These few dusty galaxies dominate the star formation within the overdensity, contributing more than the numerous Hα emitters.
Our DSFGs either lie at the upper extremes of stellar mass, dust mass, and SFR, or are entirely absent from current simulations and models. This highlights the difficulty existing models face in reproducing such extreme systems. Their properties may be driven by efficient gas inflows along cosmic filaments in protocluster environments, fueling intense star formation despite short depletion timescales (e.g. Narayanan et al. 2015).
We examine the spatial distribution of N2CLS galaxies within the overdensity, comparing their positions to the three filamentary structures identified by Sun et al. (2024). Four spectroscopically confirmed DSFGs lie along filament-1 and filament-2, while two others align with filament-3 and one may trace a new separate structure. The protocluster may be more extended than previously mapped, as several of our galaxies lie beyond current survey footprints. Confirming this will require spectroscopic redshifts for sources currently identified via photometric estimates. If all these galaxies merge by , they would form a halo significantly more massive than a typical Coma-like cluster progenitor at that epoch, emphasizing the exceptional mass scale of the environment traced by these galaxies.
The presence of such massive, dusty, and intensely star-forming galaxies at has significant implications for our understanding of galaxy formation and evolution. It suggests that the processes leading to rapid stellar and dust mass assembly were already in place within the first billion years of cosmic history in overdense environments. Future observations, particularly with facilities capable of probing the ISM properties and kinematics of these galaxies, will be crucial in unraveling the mechanisms driving their evolution.
Acknowledgements.
We acknowledge financial support from the Programme National de Cosmologie and Galaxies (PNCG) funded by CNRS/INSU-IN2P3-INP, CEA and CNES, France, and from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project CONCERTO, grant agreement No 788212). This work is based on observations carried out under project numbers 192-16 with the IRAM 30-m telescope, and projects W21CV, W23CX, and S24CF with NOEMA. IRAM is supported by INSU/CNRS (France), MPG (Germany) and IGN (Spain). This work is also based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The observations are associated with programs #1181, 1895 and 3577. Some of the data products used here were retrieved from the Dawn JWST Archive (DJA). DJA is an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140. We thank Eric F. Jiménez-Andrade for sharing their data products from the 10 GHz Karl G. Jansky Very Large Array survey of GOODS-N. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. GL thanks Joris Witstok and Thomas Henning for enlightening discussions on dust production, and warmly thanks Fengwu Sun for sharing his Python code used to generate Figure 4.References
- Akins et al. (2023) Akins, H. B., Casey, C. M., Allen, N., et al. 2023, ApJ, 956, 61
- Alcalde Pampliega et al. (2019) Alcalde Pampliega, B., Pérez-González, P. G., Barro, G., et al. 2019, ApJ, 876, 135
- Arrabal Haro et al. (2018) Arrabal Haro, P., Rodríguez Espinosa, J. M., Muñoz-Tuñón, C., et al. 2018, MNRAS, 478, 3740
- Barger et al. (2022) Barger, A. J., Cowie, L. L., Blair, A. H., & Jones, L. H. 2022, ApJ, 934, 56
- Barro et al. (2019) Barro, G., Pérez-González, P. G., Cava, A., et al. 2019, ApJS, 243, 22
- Barrufet et al. (2023) Barrufet, L., Oesch, P. A., Bouwens, R., et al. 2023, MNRAS, 522, 3926
- Battisti et al. (2020) Battisti, A. J., Cunha, E. d., Shivaei, I., & Calzetti, D. 2020, ApJ, 888, 108
- Behroozi et al. (2019) Behroozi, P., Wechsler, R. H., Hearin, A. P., & Conroy, C. 2019, MNRAS, 488, 3143
- Berta et al. (2025) Berta, S., Lagache, G., Beelen, A., et al. 2025, A&A, 696, A193
- Berta et al. (2013) Berta, S., Lutz, D., Santini, P., et al. 2013, A&A, 551, A100
- Berta et al. (2023) Berta, S., Stanley, F., Ismail, D., et al. 2023, A&A, 678, A28
- Béthermin et al. (2017) Béthermin, M., Wu, H.-Y., Lagache, G., et al. 2017, A&A, 607, A89
- Bing (2022) Bing, L. 2022, PhD thesis, thèse de doctorat dirigée par Lagache, Guilaine Physique et sciences de la matière. Astrophysique et cosmologie Aix-Marseille 2022
- Bing et al. (2023) Bing, L., Béthermin, M., Lagache, G., et al. 2023, A&A, 677, A66
- Bolzonella et al. (2000) Bolzonella, M., Miralles, J. M., & Pelló, R. 2000, A&A, 363, 476
- Boquien et al. (2019) Boquien, M., Burgarella, D., Roehlly, Y., et al. 2019, A&A, 622, A103
- Boylan-Kolchin (2023) Boylan-Kolchin, M. 2023, Nature Astronomy, 7, 731
- Brinch et al. (2024) Brinch, M., Greve, T. R., Sanders, D. B., et al. 2024, MNRAS, 527, 6591
- Brinch et al. (2025) Brinch, M., Jin, S., Gobat, R., et al. 2025, A&A, 694, A218
- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
- Burgarella et al. (2005) Burgarella, D., Buat, V., & Iglesias-Páramo, J. 2005, MNRAS, 360, 1413
- Calvi et al. (2023) Calvi, R., Castignani, G., & Dannerbauer, H. 2023, A&A, 678, A15
- Calvi et al. (2021) Calvi, R., Dannerbauer, H., Arrabal Haro, P., et al. 2021, MNRAS, 502, 4558
- Calzetti et al. (2000) Calzetti, D., Armus, L., Bohlin, R. C., et al. 2000, ApJ, 533, 682
- Carilli & Walter (2013) Carilli, C. L. & Walter, F. 2013, ARA&A, 51, 105
- Carnall et al. (2024) Carnall, A. C., Cullen, F., McLure, R. J., et al. 2024, MNRAS, 534, 325
- Carnall et al. (2018) Carnall, A. C., McLure, R. J., Dunlop, J. S., & Davé, R. 2018, MNRAS, 480, 4379
- Ceverino et al. (2024) Ceverino, D., Nakazato, Y., Yoshida, N., Klessen, R. S., & Glover, S. C. O. 2024, A&A, 689, A244
- Chabrier (2003) Chabrier, G. 2003, ApJ, 586, L133
- Charlot & Fall (2000) Charlot, S. & Fall, S. M. 2000, ApJ, 539, 718
- Chiang et al. (2013) Chiang, Y.-K., Overzier, R., & Gebhardt, K. 2013, ApJ, 779, 127
- Chiang et al. (2017) Chiang, Y.-K., Overzier, R. A., Gebhardt, K., & Henriques, B. 2017, ApJ, 844, L23
- Cowie et al. (2017) Cowie, L. L., Barger, A. J., Hsu, L. Y., et al. 2017, ApJ, 837, 139
- da Cunha et al. (2008) da Cunha, E., Charlot, S., & Elbaz, D. 2008, MNRAS, 388, 1595
- da Cunha et al. (2013) da Cunha, E., Walter, F., Decarli, R., et al. 2013, ApJ, 765, 9
- da Cunha et al. (2015) da Cunha, E., Walter, F., Smail, I. R., et al. 2015, ApJ, 806, 110
- Dekel et al. (2023) Dekel, A., Sarkar, K. C., Birnboim, Y., Mandelker, N., & Li, Z. 2023, MNRAS, 523, 3201
- Di Cesare et al. (2023) Di Cesare, C., Graziani, L., Schneider, R., et al. 2023, MNRAS, 519, 4632
- Draine (2003) Draine, B. T. 2003, ARA&A, 41, 241
- Draine et al. (2014) Draine, B. T., Aniano, G., Krause, O., et al. 2014, ApJ, 780, 172
- Dudzevičiūtė et al. (2020) Dudzevičiūtė, U., Smail, I., Swinbank, A. M., et al. 2020, MNRAS, 494, 3828
- Dunne et al. (2022) Dunne, L., Maddox, S. J., Papadopoulos, P. P., Ivison, R. J., & Gomez, H. L. 2022, MNRAS, 517, 962
- Elbaz et al. (2018) Elbaz, D., Leiton, R., Nagar, N., et al. 2018, A&A, 616, A110
- Finkelstein et al. (2015) Finkelstein, S. L., Ryan, Russell E., J., Papovich, C., et al. 2015, ApJ, 810, 71
- Fudamoto et al. (2020) Fudamoto, Y., Oesch, P. A., Faisst, A., et al. 2020, A&A, 643, A4
- Fujimoto et al. (2022) Fujimoto, S., Brammer, G. B., Watson, D., et al. 2022, Nature, 604, 261
- Fujimoto et al. (2024) Fujimoto, S., Kohno, K., Ouchi, M., et al. 2024, ApJS, 275, 36
- Gentile et al. (2024) Gentile, F., Casey, C. M., Akins, H. B., et al. 2024, ApJ, 973, L2
- Gkogkou et al. (2023) Gkogkou, A., Béthermin, M., Lagache, G., et al. 2023, A&A, 670, A16
- Gottumukkala et al. (2024) Gottumukkala, R., Barrufet, L., Oesch, P. A., et al. 2024, MNRAS, 530, 966
- Greene et al. (2024) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2024, ApJ, 964, 39
- Greve et al. (2008) Greve, T. R., Pope, A., Scott, D., et al. 2008, MNRAS, 389, 1489
- Gruppioni et al. (2020) Gruppioni, C., Béthermin, M., Loiacono, F., et al. 2020, A&A, 643, A8
- Guaita et al. (2022) Guaita, L., Aravena, M., Gurung-Lopez, S., et al. 2022, A&A, 660, A137
- Haider et al. (2016) Haider, M., Steinhauser, D., Vogelsberger, M., et al. 2016, MNRAS, 457, 3024
- Helton et al. (2024a) Helton, J. M., Sun, F., Woodrum, C., et al. 2024a, ApJ, 962, 124
- Helton et al. (2024b) Helton, J. M., Sun, F., Woodrum, C., et al. 2024b, ApJ, 974, 41
- Herard-Demanche et al. (2025) Herard-Demanche, T., Bouwens, R. J., Oesch, P. A., et al. 2025, MNRAS, 537, 788
- Hughes et al. (1998) Hughes, D. H., Serjeant, S., Dunlop, J., et al. 1998, Nature, 394, 241
- Jiang et al. (2018) Jiang, L., Wu, J., Bian, F., et al. 2018, Nature Astronomy, 2, 962
- Jiménez-Andrade et al. (2024) Jiménez-Andrade, E. F., Murphy, E. J., Momjian, E., et al. 2024, ApJ, 972, 89
- Kakimoto et al. (2024) Kakimoto, T., Tanaka, M., Onodera, M., et al. 2024, ApJ, 963, 49
- Kashino et al. (2023) Kashino, D., Lilly, S. J., Matthee, J., et al. 2023, ApJ, 950, 66
- Khusanova et al. (2021) Khusanova, Y., Bethermin, M., Le Fèvre, O., et al. 2021, A&A, 649, A152
- Kinney et al. (1996) Kinney, A. L., Calzetti, D., Bohlin, R. C., et al. 1996, ApJ, 467, 38
- Kodra et al. (2023) Kodra, D., Andrews, B. H., Newman, J. A., et al. 2023, ApJ, 942, 36
- Labbé et al. (2023) Labbé, I., van Dokkum, P., Nelson, E., et al. 2023, Nature, 616, 266
- Leitherer et al. (1999) Leitherer, C., Schaerer, D., Goldader, J. D., et al. 1999, ApJS, 123, 3
- Lemaux et al. (2022) Lemaux, B. C., Cucciati, O., Le Fèvre, O., et al. 2022, A&A, 662, A33
- Lestrade et al. (2022) Lestrade, J. F., Désert, F. X., Lagache, G., et al. 2022, A&A, 667, A23
- Li et al. (2024a) Li, Q., Conselice, C. J., Sarron, F., et al. 2024a, arXiv e-prints, arXiv:2405.17359
- Li et al. (2019) Li, Q., Narayanan, D., & Davé, R. 2019, MNRAS, 490, 1425
- Li et al. (2024b) Li, Z., Dekel, A., Sarkar, K. C., et al. 2024b, A&A, 690, A108
- Lim et al. (2024) Lim, S., Tacchella, S., Schaye, J., et al. 2024, MNRAS, 532, 4551
- Lin et al. (2025) Lin, X., Egami, E., Sun, F., et al. 2025, arXiv e-prints, arXiv:2504.08028
- Liu et al. (2018) Liu, D., Daddi, E., Dickinson, M., et al. 2018, ApJ, 853, 172
- Madau & Dickinson (2014) Madau, P. & Dickinson, M. 2014, ARA&A, 52, 415
- Matthee et al. (2024) Matthee, J., Naidu, R. P., Brammer, G., et al. 2024, ApJ, 963, 129
- McKinney et al. (2025) McKinney, J., Casey, C. M., Long, A. S., et al. 2025, ApJ, 979, 229
- McKinney et al. (2023) McKinney, J., Manning, S. M., Cooper, O. R., et al. 2023, ApJ, 956, 72
- Morishita et al. (2025) Morishita, T., Liu, Z., Stiavelli, M., et al. 2025, ApJ, 982, 153
- Morokuma-Matsui et al. (2025) Morokuma-Matsui, K., Yajima, H., & Abe, M. 2025, MNRAS, 539, 1834
- Narayanan et al. (2015) Narayanan, D., Turk, M., Feldmann, R., et al. 2015, Nature, 525, 496
- Neri et al. (2014) Neri, R., Downes, D., Cox, P., & Walter, F. 2014, A&A, 562, A35
- Noll et al. (2009) Noll, S., Burgarella, D., Giovannoli, E., et al. 2009, A&A, 507, 1793
- Oesch et al. (2023) Oesch, P. A., Brammer, G., Naidu, R. P., et al. 2023, MNRAS, 525, 2864
- Owen (2018) Owen, F. N. 2018, ApJS, 235, 34
- Pacifici et al. (2023) Pacifici, C., Iyer, K. G., Mobasher, B., et al. 2023, ApJ, 944, 141
- Perera et al. (2008) Perera, T. A., Chapin, E. L., Austermann, J. E., et al. 2008, MNRAS, 391, 1227
- Pérez-González et al. (2024) Pérez-González, P. G., Barro, G., Rieke, G. H., et al. 2024, ApJ, 968, 4
- Planck Collaboration et al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., et al. 2020, A&A, 641, A6
- Ponthieu et al. (2025) Ponthieu, N., Désert, F.-X., Beelen, A., et al. 2025, sub. to A&A
- Pope et al. (2005) Pope, A., Borys, C., Scott, D., et al. 2005, MNRAS, 358, 149
- Popesso et al. (2023) Popesso, P., Concas, A., Cresci, G., et al. 2023, MNRAS, 519, 1526
- Riechers et al. (2020) Riechers, D. A., Hodge, J. A., Pavesi, R., et al. 2020, ApJ, 895, 81
- Rodighiero et al. (2023) Rodighiero, G., Bisigello, L., Iani, E., et al. 2023, MNRAS, 518, L19
- Sommovigo et al. (2022) Sommovigo, L., Ferrara, A., Carniani, S., et al. 2022, MNRAS, 517, 5930
- Speagle et al. (2014) Speagle, J. S., Steinhardt, C. L., Capak, P. L., & Silverman, J. D. 2014, ApJS, 214, 15
- Sun et al. (2024) Sun, F., Helton, J. M., Egami, E., et al. 2024, ApJ, 961, 69
- Tacconi et al. (2020) Tacconi, L. J., Genzel, R., & Sternberg, A. 2020, ARA&A, 58, 157
- Tanaka et al. (2024) Tanaka, M., Onodera, M., Shimakawa, R., et al. 2024, ApJ, 970, 59
- Traina et al. (2024) Traina, A., Gruppioni, C., Delvecchio, I., et al. 2024, A&A, 681, A118
- Triani et al. (2020) Triani, D. P., Sinha, M., Croton, D. J., Pacifici, C., & Dwek, E. 2020, MNRAS, 493, 2490
- Walter et al. (2012) Walter, F., Decarli, R., Carilli, C., et al. 2012, Nature, 486, 233
- Wang et al. (2023) Wang, F., Yang, J., Hennawi, J. F., et al. 2023, ApJ, 951, L4
- Wang et al. (2019) Wang, T., Schreiber, C., Elbaz, D., et al. 2019, Nature, 572, 211
- Wang et al. (2024) Wang, T., Sun, H., Zhou, L., et al. 2024, arXiv e-prints, arXiv:2403.02399
- Weibel et al. (2024) Weibel, A., Oesch, P. A., Barrufet, L., et al. 2024, MNRAS, 533, 1808
- Witstok et al. (2023) Witstok, J., Jones, G. C., Maiolino, R., Smit, R., & Schneider, R. 2023, MNRAS, 523, 3119
- Xiao et al. (2024) Xiao, M., Oesch, P. A., Elbaz, D., et al. 2024, Nature, 635, 311
- Xiao et al. (2023) Xiao, M. Y., Elbaz, D., Gómez-Guijarro, C., et al. 2023, A&A, 672, A18
- Zavala et al. (2023) Zavala, J. A., Buat, V., Casey, C. M., et al. 2023, ApJ, 943, L9
- Zavala et al. (2021) Zavala, J. A., Casey, C. M., Manning, S. M., et al. 2021, ApJ, 909, 165
- Zavala et al. (2022) Zavala, J. A., Casey, C. M., Spilker, J., et al. 2022, ApJ, 933, 242
Appendix A Dust production
Observations have shown a relation between dust mass and stellar mass across an extensive redshift range. At all redshifts, the dust masses increase linearly with stellar mass. However, at fixed stellar mass, there is a large scatter of dust masses around the median, revealing the existence of a dust-poor galaxy population across all stellar masses (Triani et al. 2020). We compare on Fig. 5 our stellar-mass estimates with our inferred dust masses for the ten DSFG. We also show the stellar mass–dust mass relations derived by Witstok et al. (2023), based on dust production from stellar sources (including supernovae and AGB stars ; with and without reverse shock destruction which can lead to 95% dust destruction). Additionally, we include the relation corresponding to the highest possible metal yield (i.e. assuming all metals are locked up in dust grains), representing the theoretical maximum dust mass. Note that the exact normalization of these relations remains uncertain due to systematic effects, including the choice of IMF used to compute dust yields for a given stellar mass, and the range of plausible dust absorption cross-sections.
The dust masses observed in the majority of our galaxies (all but two) point to efficient dust formation. Indeed most of our galaxies are consistent with the dust masses expected from a simple, IMF-averaged yield of stellar sources alone (i.e., SN and AGB stars), assuming no dust destruction in the ISM or ejection from the galaxy. In general, it is challenging to populate our observed M⋆–Mdust relation using cosmological simulations, as our dusty galaxies occupy the upper envelope in both stellar and dust mass. They are, for instance, only sparsely represented in Di Cesare et al. (2023) and entirely absent in the sample of Triani et al. (2020).
The two most massive galaxies in our sample fall into regions where reverse-shock destruction of dust becomes significant, implying that additional dust growth mechanisms, such as grain growth in the interstellar medium, may be necessary to account for their observed dust masses. However, we caution that these two sources currently lack spectroscopic redshifts, and their stellar masses based on photometric estimates may be significantly overestimated. Robust spectroscopic confirmation is essential before drawing firm conclusions about their dust production.
For comparison, we also present the complete sample of N2CLS galaxies at from Berta et al. (2025), which occupy the same region i.e. near the SN+AGB dust production line. Notably, three galaxies lie above the maximum dust mass allowed: N2GN_1_17a (=3.19) and N2GN_1_32 (=3.65; detected in the Chandra 2 Ms map and requiring an AGN-torus component in the IRAC and MIPS bands) remain consistent within 1.3 (in Mdust), while N2GN_1_04 (=4.056, also known as GN20) is significantly offset.
The distribution of our measurements in the Mdust-M⋆ plane appears relatively flat. Sommovigo et al. (2022) showed that the Mdust–M⋆ relation differs between REBELS galaxies and ALPINE galaxies (even when using consistent [CII]-based methods to estimate Mdust), with REBELS galaxies exhibiting a flat trend, while ALPINE galaxies, as well as theoretical models (see Fig. 5 and their Fig. 3), suggest that more massive galaxies are dustier. Both samples probe lower dust and stellar masses than our NIKA2 sources. The observed flatness of the Mdust–M⋆ relation could partly arise from observational biases: the REBELS survey covers a limited stellar mass range, while for NIKA2, the 1.2 mm flux selection imposes a dust mass detection threshold.
Appendix B Photometric redshifts with New-Hyperz
For the five sources with literature-based photometric redshifts at 5.2 (see B25 and references therein), we re-derived their redshifts independently using the updated photometry and the New-Hyperz SED-fitting code. New-Hyperz is a modified version of the public code Hyperz originally developped by Bolzonella et al. (2000). The library of templates used for this need includes: i) Eight evolutionary synthetic SEDs computed with the Bruzual & Charlot code (Bruzual & Charlot 2003), with Chabrier IMF (Chabrier 2003) and solar metallicity, matching the observed colors of local galaxies from E to Im types (namely a delta burst Single Stellar Population, a constant star-forming system, and six -models with exponentially decaying SFR); ii) Six empirical templates for galaxies in the local Universe: four SEDs compiled by Coleman et al. (1980), and two starburst galaxies from the Kinney et al. 1996 library; iii) Ten Starbursts99 evolutionary models (Leitherer et al. 1999), including e-lines, for single bursts and constant SFR models, each spanning five metallicities (Z 0.04, 0.02, 0.008, 0.004 and 0.001). The explored redshift domain goes from z0 to 16, without any prior in luminosity. The only restriction is that the age of the stellar population should be smaller than the age of the universe for a given redshift. Internal dust extinction is a free parameter that is allowed to take values between A0 and 6 magnitudes following the Calzetti et al. (2000) extinction law. During the SED-fitting process, to account in a realistic way for calibration uncertainties related to the combination of data coming from different facilities, across a wide range in wavelengths, we imposed a minimum photometric error of 0.2 magnitudes which is quadratically added to the original error bars. Indeed, in our experience this is needed to prevent artificially small error bars leading to low-quality fits, by correcting from effects such as biases between different instruments due to (absolute) zero-point uncertainties, photometry computed with an effectively different spatial sampling and spatial resolution, and contribution to the total flux coming from different spatial regions.
The quality of the photometric redshifts with New-Hyperz and the settings described above for the SED-fitting has been checked using the 29 N2CLS galaxies with spectroscopic redshifts (Sect. 2), excluding N2GN_1_61 because it is only detected in one filter. Spectroscopic redshifts are in the range between z0.6 and 7.2. The results are as follows, with : = 0.058, the median of = -0.016, and the normalized median absolute deviation, which is less sensitive to outliers: = 1.48 median = 0.018. Regarding the fraction of outliers, defined in a conservative way as sources with , it is found to be 21% (6 objects/28), and only two galaxies (7%) exhibit . Among these six outliers, four exhibit a degenerate solution with a secondary solution consistent with . One of them is an X-ray AGN at zspec = 3.65 with a NIR-MIR SED dominated by a torus component, exhibiting a degenerate solution between z and z. Note that a poor zphot estimate is obtained for an obscured quasar at zspec=7.19 (Fujimoto et al. 2022).
Appendix C Contribution to the cosmic SFRD
Based on the five N2CLS galaxies with spectroscopic redshifts, we derive a cosmic star formation rate density (SFRD) of M⊙ yr-1 Mpc-3. Including the additional contribution from sources with photometric redshifts raises the total SFRD to M⊙ yr-1 Mpc-3. The quoted uncertainties combine Poisson errors and uncertainties on individual SFR estimates in quadrature. Figure 6 shows the contribution of the N2CLS galaxies within the overdensity at z to the cosmic SFRD. For comparison, we also display the total SFRD corrected for dust attenuation as compiled by Madau & Dickinson (2014). This comparison is discussed in the main text (Sect. 7.1). Additionally, we include the dust-obscured and unobscured components of the SFRD from Zavala et al. (2021). In their analysis, the contribution from dust-obscured galaxies peaks at –3 and declines at higher redshifts, reaching levels comparable to the unobscured star formation traced by rest-frame UV/optical surveys by . We find that the contribution of the dusty galaxies in the z overdensity is more than five times higher than the dust-obscured SFRD reported by Zavala et al. (2021) at similar redshifts.
Appendix D Notes about individual sources and JWST cutouts
Below, we provide a detailed description of each individual source identified at z5.2 (but HDF850.1, GN10 and N2GN_1_13). The JWST photometry measurements follow a similar approach to that described by Weibel et al. (2024), with some modifications as outlined in Section 2.8 of B25. In Figure 7, we present the multi-wavelength JWST postage stamps, where available.
N2GN_1_20 is detected by SMA. SMA position lies at 1.2 arcsec from S-CANDELS J123550.35+621041.8. It has a Ks band flux of 0.75 Jy, resulting in a very red Ks to 3.6 m color, , interpreted as a deep break at 4000 Å combined to a large extinction, that leads to a (Liu et al. 2018). Using New-Hyperz, we obtained =5.16, the broad error bars being the result of a degenerated solution in the parameter space between redshit, extinction and age of the stellar population. This source is not covered by JWST.
N2GN_1_23: is detected by JWST/FRESCO at =5.179 and is the so-called S3 source extensively discussed in Xiao et al. (2024). We have a firm identification of our N2CLS source through our NOEMA follow-up program (see B25) with the NOEMA counterpart lying at a distance of 0.167 arcsec. The NIKA2 source has both a SCUBA2 450 and 850 m counterparts but the NOEMA position lies at 3.6 arcsec from the 450 m source Number 78 of Barger et al. (2022), identified with a galaxy at z=0.29 (see their Table 2). The 850 m source is identified with SMM123656621205 (Number 80), which is at 2.8 arcsec from the NOEMA position, and with a redshift estimate of 3.75 (Cowie et al. 2017). The high-z source is clearly blended with a nearby object and flux measurements from Herschel and SCUBA2 450 m cannot be used to construct the multi-wavelength SED. The NOEMA source has both a 1.4 GHz (Owen 2018) and 10 GHz (Jiménez-Andrade et al. 2024) radio counterpart.
N2GN_1_26 is identified with a high- HST galaxy thanks to SMA data. Kodra et al. (2023) give a photometric redshift . The FIR SED is well defined by the PACS, SCUBA2, SMA, and NIKA2 data. The 24 m flux is well matched by an intense 3.3 m PAH feature. There is a 10 GHz source at the HST position. This source is newly observed and detected by the PANORAMIC survey with F277W, F356W. Using New-Hyperz, it has a photometric redshift of 4.87, consistent with previous findings at 1 level.
N2GN_1_33: corresponds to the SMA GOODS-N source number 13 (Cowie et al. 2017). Three sources lie inside the SMA beam: a bright optical galaxy, a very faint optical galaxy revealed by JWST, and a radio source which lies in the middle of the two optical sources. The bright and faint optical galaxies are respectively at 1.21 and 0.56 arcsec from the SMA position. We chose the closest one as the identification. This is confirmed by the 10 GHz low- and high-resolution maps which have a clear source at this position. In B25, it has ; using New-Hyperz, we obtain , in perfect agreement with previous findings.
N2GN_1_43 is identified with NOEMA. It is a K-band dark galaxy. It is detected by JWST at in the F277W band, in F200W, and tentatively by HST at shorter wavelengths. Although not in the IRAC catalog, it is clearly seen in the maps at 4.5, 5.8 and 8 m. It has also a clear 10 GHz signal (Jiménez-Andrade et al. 2024) at the JWST source position. A fit with CIGALE in B25 places it at with a 4000 Å break between the two JWST measurements mentioned here. The photometric redshift from New-Hyperz is 5.31, consistent with previous estimates, with a degenerated solution in the parameter space between redshift and extinction leading to broad error bars.
N2GN_1_57: The VLA source, coinciding with GN4 26261 (MIOP program) at z=0.9, is likely not the right counterpart as it lies at 4.9 arcsec from the N2CLS position. Closer to the N2CLS position (2.2 arcsec), the source HRG14 J123731.66+621616.7 (Finkelstein et al. 2015) (also detected by the SHARDS middle-band survey) is at a photometric redshift of z=5.20.06 in Arrabal Haro et al. (2018) and z=5.30 in Kodra et al. (2023). It is detected in the four IRAC channels and has a pronounced 4000 Å break at m. The photometric redshift obtained with New-Hyperz is z=5.18, consistent with previous findings.
N2GN_1_61 is not only optically-dark but also IRAC-dark ([4.5]¿24.6 mag at 5). The accurate position is given by NOEMA (see B25) revealing a single source with a flux of 69884Jy at 1.18 mm, in perfect agreement with the NIKA2 flux. JWST grism observations from FRESCO detected one emission line, consistent with H at (M. Xiao, priv. comm.). It is observed by JWST with F090W, F115W, F182M, F210M, F356W, F444W filters, but only detected by F356W and F444W (see Fig. 7). The observed SED consists of only three data points (NIKA2 1.2 mm, SCUBA2 850 m and JWST F444W), but the availability of the redshift allows us to perform a fit. Derived parameters have large uncertainties (but are consistent between MAGPHYS-highz and CIGALE).